# Conformal Validity Guarantees Exist for Any Data Distribution

### v2

@@ 4.3 Deriving Weighted CP Validity Guarantees and Algorithms for MFCS, or Any Data Distribution
A General Procedure for Deriving Weighted CP Algorithms with Valid Coverage Guarantees To improve the accessibility and practical utility of our paper’s main theoretical contribution, we now outline a general procedure that can be used to derive weighted CP algorithms for any specific joint PDF $f$ over the calibration and test data. Theorem 4.1 implies the existence of this procedure without premising it on any assumptions on $f$, while emphasizing that the only role of distributional assumptions like exchangeability is for tractability in the last step.
@@
    List assumptions (if any): For a given application setting, specify assumptions on $f$, if any, such as conditional independence or invariance assumptions. (For example, the MFCS assumptions are formally given by Eq. (3); informally, MFCS assumes $Y|X$ is invariant, while $X$ is dynamic depending on past observations.) Alternatively, one could begin by describing an application’s data-generating process by a probablistic graphical model (e.g., a Markov random field, Bayesian network, or a causal directed acyclic graph) and use the conditional independence assumptions implied by that graphical model (for introductions to such models see Koller & Friedman (2009); Pearl (2009)). [^7] Probablistic graphical models can intuitively represent a data-generating process as a graph with nodes corresponding to random variables (or random vectors). Missing edges in such a graph imply conditional independences in the joint PDF $f$; the converse depends on further assumptions (Pearl, 2009).
@@
Deriving CP for Agent-Induced MFCS For a concrete demonstration of how this general procedure can be used to derive CP methods for a given practical setting, we now return to the MFCS setting introduced in Section 2.3, which characterizes common instances of ML-agent-induced feedback-loop data shifts. We provide only a sketch of key ideas here; for full details see Appendix B. Step 1 of this procedure for MFCS begins by recalling Eq. (3), the conditional independence or invariance assumptions that we used to define MFCS. Using standard probability rules and leveraging these MFCS assumptions, for Step 2 we are able to factorize the joint PDF $f$ to separate the time-dependent factors corresponding to $X_{j}\mid Z_{1},…,Z_{j-1}$ from the time-invariant factors corresponding to $Y\mid X$:
@@
$\prod_{j=1}^{n+t}\Big{[}\underbrace{\mathbb{P}(x_{j}\ |\ z_{1},...,z_{j-1})}_{\text{Time-dependent factors}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(y_{j}|x_{j})\Big{]}}_{\text{Time-invariant factor}}.$ (7)
@@
$=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$ (8)
after canceling out $\prod_{j=1}^{n+t}\big{[}\mathbb{P}(y_{\sigma(j)}|x_{\sigma(j)})\big{]}$ from the numerator and denominator, as it is invariant to permutations $\sigma$.
As a corollary to Theorem 4.1, a CP method with weights given by Eq. (4.3) has a valid coverage guarantee premised on the MFCS assumptions in Eq. (3). We can thus turn to practical estimation. Firstly, in the agent-induced MFCS settings that we focus on in our experimental evaluations (Section 5), the epistemic challenge is overcome because the dynamic components of $f$ are known entirely. That is, the invariant factor corresponding to $Y|X$ cancels in the ratio, and in our practical evaluations the remaining factors $\mathbb{P}(x|Z_{1},...,Z_{j-1})$ represent ML-agent-controlled query probability functions at each timestep $j$. Next, we examine computational complexity. Due to the MFCS IID initialization implying $\mathbb{P}(x|Z_{1},...,Z_{j-1})=\mathbb{P}(x)$ for $j\leq n$, and when $\mathbb{P}(x|Z_{1},...,Z_{j-1})$ is computed from an ML model that treats $Z_{1},...,Z_{j-1}$ symmetrically (as in our experiments), the exact computation of Eq. (4.3) has complexity $\mathcal{O}(\prod_{j=1}^{t}(n+j))$. Though reduced from $\mathcal{O}((n+t)!)$, the complexity for arbitrary $f$, this this still quickly becomes intractable for large $t$. To alleviate this remaining complexity bottleneck, we propose estimating Eq. (4.3) by using only the “highest-order” terms from the $d$ most recent timesteps. That is, for a user-specified estimation depth $d\in[t]$, then we define our “$d$-step” estimate of Eq. (4.3) as
@@
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{j=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n+t+1-d}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod^{n+t}_{j=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n+t+1-d}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$
@@
Figure 1: Results for fluorescent protein design ($n=32$, $\lambda=8$) with a linear ridge regressor, comparing the proposed MFCS full CP methods to standard full CP and one-step weighted full CP baselines. We repeated this experiment with 1,000 random seeds; the error bars for mean coverage and mean predicted fitness represent standard error, while the error bars for median interval width represent interquartile range. The proposed MFCS full CP maintains coverage where the baselines do not. Its intervals are wider than those of the one-step FCS baseline where the latter loses coverage ($t=2,3,4$), but similar where that baseline maintains coverage ($t=1,5$).
@@ 5 Experimental Results
We evaluate the practical performance of our proposed CP methods for ML-agent-induced MFCS on synthetic black-box optimization (specifically protein design) and active learning tasks. All experiments focus on a regression setting with all CP methods using the absolute residual score $\widehat{\mathcal{S}}(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}_{t}(x)$ is the predicted mean. Our proposed methods are defined by estimating Eq. (6) with weights as in Eq. (9). The primary desired criterion for evaluating a method’s reliability is whether its prediction intervals maintain empirical coverage at or above the target coverage level $1-\alpha$. Secondarily—if target coverage is achieved—then smaller predictive sets are more informative (i.e., predicting the full label space $\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$ could trigger deferral to an expert but does not itself convey new insights from the predictor). In all experiments we compare to Exchangeable CP (Vovk et al., 2005) and One-Step FCS CP (Fannjiang et al., 2022), and in the active learning experiments we additionally compare with Adaptive Conformal Inference (ACI) (Gibbs & Candès, 2021). Our proposal distributions follow the form $\mathbb{P}_{t}(X_{n+t}=x):=\mathbb{P}(X_{n+t}=x\mid Z_{1:(n+t-1)})\propto\exp(\lambda\cdot u_{t}(x))$, where $u_{t}$ is some utility function and $\lambda\geq 0$ is a hyperparameter that controls the magnitude of agent-induced covariate shift. In all experiments we corrupt observed labels with Gaussian noise to simulate noisy observations. We add every query point to the training set for the full CP experiments, and for split CP we randomly assign the query points to either training or calibration by flipping a fair coin. (Details in Appendix D.)
@@ 5.1 Multistep Protein Design Experiments
Optimization-Induced MFCS Experimental Details Our protein design procedure is similar to a common approach in the biomolecular engineering literature (Biswas et al., 2021; Zhu et al., 2024) and related works in CP (Fannjiang et al., 2022; Prinster et al., 2023), but extended to a more realistic online setting where protein sequence proposals and model updates are interleaved and dependent (Stanton et al., 2022; Gruver et al., 2023; Angermueller et al., 2023). All methods start with an IID initialization by sampling $n$ points uniformly at random from a combinatorially complete library (Poelwijk et al., 2019), splitting the data equally into training and calibration sets if using split CP. See Appendix D.1 for additional details. We evaluate the mean coverage and median intervals width as a function of the number of optimization steps past initialization, fitting the regressor to $Z_{1:n+t-1}$, and sampling one query point with replacement (with added measurement noise) from the library with probability $\mathbb{P}_{t}(X_{n+t}=x)\propto\exp(\lambda\cdot\widehat{\mu}_{t}(x))$.
Figure 2: Results for fluorescent protein design with an MLP regressor ($n=64$, $\lambda=5$), comparing the proposed MFCS Split CP methods to standard split CP and one-step weighted split CP baselines. The experiment is repeated with 500 random seeds; the error bars for mean coverage and mean predicted fitness represent standard error, while for median interval width they represent the interquartile range. Error bars extending beyond the top of the figure indicate that the upper quartile is infinite. MFCS Split CP maintains coverage, but its interval are occasionally very wide, suggesting the proposal distribution is too aggressive for consistently informative uncertainty quantification.
Full CP Results for Multistep Protein Design Our first evaluations focus on variants of Full CP (with leave-one-out scores; see Appendix A.2), which efficiently uses the same dataset for both training and calibration (often resulting in more accurate and sharper prediction intervals relative to other CP methods). Full CP’s data efficiency comes at a steep computational cost, with its complexity corresponding to the number of predictors that must be trained: For an arbitrary $\widehat{\mu}$, computing the MFCS Full CP weights in Eq. (4.3) requires training $|\mathcal{Y}|\cdot\prod_{j=1}^{t}(n+j)$ distinct predictors, where $|\mathcal{Y}|$ is the cardinality of the label space. [^8] In practice, regression labels are discretized by a fine grid. If $\widehat{\mu}$ is a ridge regression predictor, for the training cost is reduced to $\prod_{j=1}^{t}(n+j)$, and we moreover set our estimation depth to $d=2$ for computing our estimated MFCS weights Eq. (9), which reduces the training cost to $(n+t)\cdot(n+t-1)$.
@@
Figure 1 reports the results for our proposed MFCS Full CP method (blue circles) with hyperparameters reported in the caption. On both the blue and red fluorescent protein design datasets, and on all five evaluated design steps, our proposed MFCS Full CP method maintains coverage above the target level even when both baselines lose coverage below this level. The MFCS method’s coverage is maintained by increasing the conservativeness (size) of its prediction intervals relative to the One-Step baseline after the first design step. The superimposed fitness curves on the right verify that differences in coverage can be attributed to the different CP methods, not the data being collected.
@@
Figure 3: Active learning experiments of proposed multistep split CP method for $d=3$ (red) compared to baselines of unweighted split CP (orange), one-step split CP (green), and ACI (gray). The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. For all experiments, a holdout test set of 250 samples was first sampled uniformly at random from the full dataset to track the accuracy of the base predictor as the active learning progressed. To simulate a practical active learning setting where there often tends to be sample-selection bias in how the initial training data is obtained relative to the desired test distribution, we sampled the initial training and calibration datasets from the full dataset with probabilities proportional to $\exp(\gamma\cdot X_{\text{PCA}1}^{\text{normed}})$, where $X_{\text{PCA}1}^{\text{normed}}$ is the min-max-normed first principal component representation of the data. All values are computed over 350 distinct random seeds.
Efficient Computation with Split CP for Multistep Protein Design While the training cost of Full CP is prohibitive for many types of ML predictors, this limitation is avoided by variants of Split CP (Papadopoulos, 2008), which we also extend to MFCS. Split CP methods reduce computational effort by splitting the available data into a “proper” training set used only to fit a single ML model, and a separate holdout “calibration” set on which the fitted model computes CP scores to construct the CP set. Split CP’s computational complexity thus corresponds to the evaluation cost of the pre-fitted model on the calibration and test data.
@@
The results in Figure 2 use a neural network regressor on the protein design datasets to evaluate our MFCS Split CP method. We moreover compare the effect of several estimation depths $d$ for estimating the MFCS weights as in Eq. (9). For weight estimation depths $d\in\{3,4\}$, our proposed MFCS CP methods maintain target coverage across all 10 evaluated design steps, for both the blue and red fluorescent protein design datasets, even when the baselines break down. Increasing the estimation depth tends to improve coverage at later design steps, for instance relative to the one-step FCS, apparently with diminishing returns (the red 3-step and purple 4-step methods have nearly overlapping performance). Coverage is maintained by the MFCS Split CP methods by slightly increasing the size of the intervals relative to the One-Step baseline at later design steps. However, both the MFCS Split CP method and the One-step FCS Split CP method occasionally have noninformative (infinite width) intervals, which could be unacceptable in some application settings. In Appendix E, we show that adaptively bounding the ML agent’s query probability function can achieve consistently informative (finite) interval widths without losing coverage, by restricting the ML agent’s feedback shift.
5.2 Active Learning Induced MFCS

### v3

@@ 4.3 How to Find Weighted CP Validity Guarantees and Algorithms for MFCS (or Any Data Distribution)
A General Procedure for Deriving Weighted CP Algorithms with Valid Coverage Guarantees To improve the accessibility and practical utility of our paper’s main theoretical contribution, we now outline a general procedure that can be used to derive weighted CP algorithms for any specific joint PDF (or mass function) $f$ over the calibration and test data. Theorem 4.1 implies the existence of this procedure without premising it on any assumptions on $f$, while emphasizing that the only role of distributional assumptions like exchangeability is for tractability in the last step.
@@
    List assumptions (if any): For a given application setting, specify assumptions on $f$, if any, such as conditional independence or invariance assumptions. (For example, the MFCS assumptions are formally given by Eq. (3); informally, MFCS assumes $Y|X$ is invariant, while $X$ is dynamic depending on past observations.) Alternatively, one could begin by describing an application’s data-generating process by a probablistic graphical model (e.g., a Markov random field or a causal directed acyclic graph) and use the conditional independence assumptions implied by that graphical model (for an example see Appendix B.3; for details on such models Koller & Friedman (2009); Pearl (2009)). [^7] Probablistic graphical models can intuitively represent a data-generating process as a graph with nodes corresponding to random variables (or random vectors). Missing edges in such a graph imply conditional independences in the joint PDF $f$; the converse depends on further assumptions (Pearl, 2009).
@@
Deriving CP for Agent-Induced MFCS For a concrete demonstration of how this general procedure can be used to derive weighted CP methods for a given practical setting, we now return to the MFCS setting introduced in Section 2.3, which characterizes common instances of ML-agent-induced feedback-loop data shifts. We provide only a sketch of key ideas here; for full details see Appendix B. Step 1 of this procedure for MFCS begins by recalling Eq. (3), the conditional independence and invariance assumptions that we used to define MFCS. Using standard probability rules and leveraging these MFCS assumptions, for Step 2 we are able to factorize the joint PDF $f$ to separate the time-dependent factors corresponding to $X_{j}\mid Z_{1},…,Z_{j-1}$ from the time-invariant factors corresponding to $Y\mid X$:
@@
$\prod_{j=1}^{n+t}\Big{[}\underbrace{p(x_{j}\ |\ z_{1},...,z_{j-1})}_{\text{Time-dependent factors}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}p(y_{j}|x_{j})\Big{]}}_{\text{Time-invariant factor}}.$ (7)
@@
$=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}p(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}p(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$ (8)
after canceling out $\prod_{j=1}^{n+t}p(y_{\sigma(j)}|x_{\sigma(j)})$ from the numerator and denominator, as it is invariant to permutations $\sigma$.
As a corollary to Theorem 4.1, a CP method with weights given by Eq. (4.3) has a valid coverage guarantee premised on the MFCS assumptions in Eq. (3). We can thus turn to practical estimation. Firstly, in the agent-induced MFCS settings that we focus on in our experimental evaluations (Section 5), the epistemic challenge is overcome because the dynamic components of $f$ are known entirely. That is, the invariant factor corresponding to $Y|X$ cancels in the ratio, and in our practical evaluations the remaining factors $p(x|Z_{1},...,Z_{j-1})$ represent ML-agent-controlled query probability functions at each timestep $j$. Next, we examine computational complexity. Due to the MFCS IID initialization implying $p(x|Z_{1},...,Z_{j-1})=p(x)$ for $j\leq n$, and when $p(x|Z_{1},...,Z_{j-1})$ is computed from an ML model that treats $Z_{1},...,Z_{j-1}$ symmetrically (as in our experiments), the exact computation of Eq. (4.3) has complexity $\mathcal{O}(\prod_{j=1}^{t}(n+j))$. Though reduced from $\mathcal{O}((n+t)!)$, the complexity for arbitrary $f$, this still quickly becomes intractable for large $t$. To alleviate this remaining complexity bottleneck, we propose estimating Eq. (4.3) by using only the “highest-order” terms from the $d$ most recent timesteps. That is, for a user-specified estimation depth $d\in\{1,...,t\}$, we define our “$d$-step” estimate of Eq. (4.3) as
@@
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{j=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n+t+1-d}p(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod^{n+t}_{j=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n+t+1-d}p(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$
@@
Figure 1: Results for protein design ($n=32$, $\lambda=8$) with linear ridge regression, comparing the proposed MFCS Full CP method ($d=2$) to standard Full CP and One-Step FCS Full CP. Values are computed over 1,000 random seeds; error bars for mean coverage and mean predicted fitness are standard errors, while error bars for median interval width are interquartile ranges. MFCS Full CP maintains coverage where the baselines do not. Its intervals are wider than those of the one-step FCS baseline where the latter loses coverage ($t=2,3,4$), but similar where the baseline maintains coverage ($t=1,5$). (Further experimental details in Appendix E.1 Table 1.)
Full versus Split CP Algorithms The derivation procedure and MFCS methods presented in this section apply to variants of both Full and Split CP. Full CP efficiently uses the same dataset for both training and calibration (often resulting in more accurate and sharper prediction intervals). Full CP’s data efficiency comes at a steep computational cost, with its complexity dominated by the number of predictors that must be trained for each candidate label: For an arbitrary predictor, computing the (leave-one-out) Full CP MFCS weights in Eq. (4.3) requires training $|\mathcal{Y}|\cdot\prod_{j=1}^{t}(n+j)$ distinct predictors, where $|\mathcal{Y}|$ is the cardinality of the label space. [^8] In practice for regression, a grid approximates the label space. Certain predictor classes (e.g., linear ridge regression) have closed-form solutions as a function of $y$ (Vovk et al., 2005), which reduces the training cost to $\prod_{j=1}^{t}(n+j)$.
Variants of Split CP (Papadopoulos, 2008) avoid the training cost of Full CP by splitting the available data into a “proper” training set used only to fit a single ML model (for each updated training dataset), and a separate holdout “calibration” set on which the fitted model computes CP scores to construct the CP set. Split CP’s computational complexity at inference time thus corresponds to the evaluation cost of the pre-fitted model on the calibration and test data.
Figure 2: Results for fluorescent protein design with a multi-layer perceptron (MLP) regressor (initial training and calibration data sizes: $n_{\text{train}}=n_{\text{cal}}=32$; $\lambda=5$), comparing the proposed MFCS Split CP methods to Standard Split CP and the One-Step FCS Split CP baselines. Values are computed over 500 random seeds. Error bars for mean coverage and mean predicted fitness are standard errors; for median interval width they are the interquartile ranges. Error bars extending beyond the top of the figure indicate infinite upper quartiles. MFCS Split CP maintains coverage where the baselines do not, but its intervals are occasionally very wide, suggesting the proposal distribution is too aggressive for dependably informative uncertainty estimation. (More experimental details in Appendix E.1 Table 2.)
@@ 5 Experimental Results
We evaluate the performance of our proposed CP methods for ML-agent-induced MFCS on synthetic black-box optimization (specifically protein design) and active learning tasks. All experiments are a regression setting with all CP methods using the absolute residual score $\widehat{\mathcal{S}}(x,y)=|y-\widehat{\mu}_{t}(x)|$, where $\widehat{\mu}_{t}(x)$ is the prediction of the model fit to $Z_{\text{train}}^{(t)}$, the training data available just prior to time $t$ (for Split CP these are separate from calibration data, for Full CP $Z_{\text{train}}^{(t)}=Z_{1:(n+t-1)}$). Our proposed methods are defined by estimating Eq. (6) with weights as in Eq. (9). The primary desired criterion for reliability is whether its prediction intervals maintain empirical coverage at or above the target coverage level $1-\alpha$. Secondarily—if target coverage is achieved—then smaller predictive sets are more informative (i.e., predicting the full label space $\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$ could trigger deferral to an expert but does not itself convey new insights from the predictor). In all experiments we compare to Exchangeable CP (Vovk et al., 2005) and One-Step FCS CP (Fannjiang et al., 2022), and in the active learning experiments we also compare with Adaptive Conformal Inference (ACI) (Gibbs & Candès, 2021). Our proposal distributions follow the form $p(x\mid Z_{\text{train}}^{(t)})\propto\exp(\lambda\cdot u_{t}(x))$, where $u_{t}$ is some utility function and $\lambda\geq 0$ is a hyperparameter that controls the magnitude of agent-induced covariate shift. In all experiments we corrupt labels with Gaussian noise to simulate noisy observations. Each query point is added to training for full CP experiments, and for split CP, to training or calibration by flipping a fair coin. (See Appendix E for details, Appendices C and F for additional experiments.)
@@ 5.1 Multistep Protein Design Experiments
Optimization-Induced MFCS Experimental Details Our protein design procedure is similar to a common approach in the biomolecular engineering literature (Biswas et al., 2021; Zhu et al., 2024) and related works in CP (Fannjiang et al., 2022; Prinster et al., 2023), but extended to a more realistic online setting where protein sequence proposals and model updates are interleaved and dependent (Stanton et al., 2022; Gruver et al., 2023; Angermueller et al., 2023). All methods start with an IID initialization by sampling $n$ points uniformly at random from a combinatorially complete library (Poelwijk et al., 2019), splitting the data equally into training and calibration sets if using split CP. (See Appendix E.1 for details.) We evaluate the mean coverage and median intervals width as a function of $t$ (the number of optimization steps past initialization), fitting the regressor to training data observed prior to time $t$, and sampling one query point with replacement (with added measurement noise) from the library, each with probability $p(x\mid Z_{\text{train}}^{(t)})\propto\exp(\lambda\cdot\widehat{\mu}_{t}(x))$.
@@
Full CP Results Figure 1 reports the results for our proposed MFCS Full CP method (blue circles) with hyperparameters reported in the caption. On both the blue and red fluorescent protein design datasets, and on all five evaluated design steps, our proposed MFCS Full CP method maintains coverage above the target level even when both baselines lose coverage below this level. The MFCS method’s coverage is maintained by increasing the conservativeness (size) of its prediction intervals relative to the One-Step baseline after the first design step. The superimposed fitness curves on the right verify that differences in coverage can be attributed to the different CP methods, not the data collected.
@@
Figure 3: Active learning experiments of proposed MFCS Split CP method for $d=3$ (red) compared to baselines of unweighted Split CP (orange), One-Step Split CP (green), and ACI (gray). Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set to track the accuracy of the base predictor. X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct randomm seeds (full experimental details in Appendix E.2). The proposed MFCS CP method maintains target coverage over a long time horizon even where baselines do not.
@@
Split CP Results with Neural Network Regressor The results in Figure 2 use an MLP regressor on the protein design datasets to evaluate our MFCS Split CP method. We moreover compare the effect of several estimation depths $d$ for estimating the MFCS weights as in Eq. (9). For weight estimation depths $d\in\{3,4\}$, our proposed MFCS CP methods maintain target coverage across all 10 evaluated design steps, for both the blue and red fluorescent protein design datasets, even when the baselines break down. Increasing the estimation depth $d$ tends to improve coverage at later design steps, for instance relative to the One-Step FCS method, apparently with diminishing returns (the red 3-step and purple 4-step methods have nearly overlapping performance). Coverage is maintained by the MFCS Split CP methods by slightly increasing the size of the intervals relative to the one-step baseline at later design steps. However, both the MFCS Split CP method and the One-Step FCS Split CP method occasionally have noninformative (infinite width) intervals, which could be unacceptable in some application settings. In Appendix C, we show that adaptively bounding the ML agent’s query probability function can achieve dependably informative (finite) interval widths without losing coverage, by restricting the ML agent’s feedback shift.
5.2 Active Learning-Induced MFCS